# Ando type dilation for completely contractive covariant representations

### v1

@@ Abstract.
Solel proved that every completely contractive covariant representation of a product system $\mathbb{E}$ over $\mathbb{N}^{2}_{0}$ can be dilated to an isometric representation of $\mathbb{N}^{2}_{0},$ which covers Ando’s dilation result when $C^{*}$-algebra $\mathcal{M}=\mathbb{C}$ and $\mathbb{E}(\mathbf{n})=\mathbb{C},$ where $\mathbf{n}\in\mathbb{N}^{2}_{0}.$ That arises a natural question that exactly which type of representations dilate completely contractive covariant representation of product system $\mathbb{E}$ over $\mathbb{N}^{2}_{0}?$ To identify the isometric covariant representations for which the above has an affirmative answer is the main object of this article.
@@ 1. Introduction
The theory of dilation is a fundamental result established by Sz.-Nagy [14], which asserts that any contraction on a Hilbert space can be dilated to an isometry acting on a larger Hilbert space. This result laid the foundation for a deeper understanding of operator theory, providing a powerful tool for studying contractions through their isometric extensions. Building on Sz.-Nagy’s foundational work, Ando [1] significantly advanced the theory by constructing isometric dilations for pairs of commuting contractions. That is, for any pair of commuting contractions $(T_{1},T_{2})$ on a Hilbert space $\mathcal{H},$ then there exists a pair of commuting isometries $(V_{1},V_{2})$ on a Hilbert space $\mathcal{K}\supseteq\mathcal{H}$ such that
@@
where $P_{\mathcal{H}}$ denotes the orthogonal projection onto ${\mathcal{H}}.$ Ando’s theorem thus confirms that every pair of commuting contractions can be dilated to a pair of commuting isometries, providing a complete solution in the two-variable case. Parrott [15] gave a counterexample showing that for $n\geq 3,$ $n$-tuples of commuting contractions do not generally admit isometric dilations. This striking limitation highlights the intrinsic difficulties of multivariable dilation theory and motivates a more nuanced investigation into the specific conditions under which such dilations might still be possible.
In recent years, a growing body of literature has explored the interplay between dilation and positivity, revealing that certain classes of operator tuples under specific positivity assumptions can indeed possess isometric dilations. See [2, 4, 5, 8, 13] for more details in the polydisc setup. The main motivation behind our interest comes from a recent work by Das and Sarkar [5], they obtained an explicit isometric dilation of pair of commuting contractions in the following theorem:
@@
Muhly and Solel [12] presented isometric dilation for completely contractive covariant representations of $C^{*}$-correspondences. Solel [24] proved that every completely contractive representations of product system of $C^{*}$-correspondences over the semigroup $\mathbb{N}_{0}^{2}$ can be dilated to isometric representations. Solel [25] studied completely contractive representations of product system of $C^{*}$-correspondences over $\mathbb{N}_{0}^{k},$ and gave a necessary and sufficient condition for such a representation to have a regular isometric dilation. In [22] Skalski established some sufficient conditions for the existence of a not necessarily regular isometric dilation of a completely contractive representation of the product system and discussed the difference between regular and *-regular dilations. He also analysed the case of product systems associated with higher-rank graphs. The main purpose of this paper is to follow this direction along [5] and to investigate which type of representation dilates completely contractive covariant representations of product system.
@@ 1.1. Notations and Preliminaries
We begin this section by recalling few concepts and their properties about $C^{*}$-correspondences from [10, 12, 16, 19, 21]. Let $E$ be a Hilbert $C^{*}$-module over a $C^{*}$-algebra $\mathcal{M}.$ By $B^{a}(E)$ we denote the $C^{*}$-algebra of all adjointable operators on $E$. We say that the module $E$ is a $C^{*}$-correspondence over $\mathcal{M}$ if there exists a left $\mathcal{M}$-module structure through a non-zero *-homomorphism $\phi:\mathcal{M}\to B^{a}(E)$ in the following sense
@@
for every $\eta_{1},\eta_{2}\in F,$ $\xi_{1},\xi_{2}\in E$ and $a\in\mathcal{M}.$ In this paper, we adopt the following notations: $\mathcal{H}$ denotes a Hilbert space, $E$ represents a $C^{*}$-correspondence over $\mathcal{M}$, and $B(\mathcal{H})$ refers to the algebra of all bounded linear operators on $\mathcal{H}.$ The notion of completely contractive covariant representations plays a crucial role in this paper.
@@
Suppose that $(\sigma,T)$ is a completely bounded covariant representation of ${E}$ on $\mathcal{H}.$ A closed subspace $\mathcal{W}$ of $\mathcal{H}$ is said to be $(\sigma,T)$-invariant $(resp.(\sigma,T)$-reducing) (cf. [21]) if it is $\sigma(\mathcal{M})$-invariant and (resp. both $\mathcal{W},\mathcal{W}^{\perp}$) is invariant by each operator $T(\xi)$ for all $\xi\in E.$ The restriction provides a new representation $(\sigma,T)|_{\mathcal{W}}$ of $E$ on $\mathcal{W}.$
@@
The following theorem is an analogue of [14, Theorem 3.2].
@@
$\mathcal{H}_{0}:$ =$=\{h\in\mathcal{H}:\|(I_{E^{\otimes n}}\otimes\widetilde{T}_{m})(\xi_{n+m}\otimes h)\|=\|\xi_{n+m}\otimes h\|\quad\mbox{and}\quad$
@@ 1.2. Completely contractive representations over product system
Suppose $u=\{u_{1},u_{2},\dots,u_{r}\}\subseteq\{1,2,\dots,k\}$ such that $u_{1}<u_{2}<\dots<u_{r}.$ Define $\mathbf{e}(u):=\mathbf{e}_{u_{1}}+\mathbf{e}_{u_{2}}+\dots+\mathbf{e}_{u_{r}},$ where $\mathbf{e}_{i}:=(0,0,\dots,1,0,\dots 0),$ i.e., $1$ at $i^{th}$ place and 00 otherwise. Next we define the following Brehmer-Solel condition:
@@
The following theorem gave an isometric dilation of completely contractive covariant representations of product system due to Rohilla, Trivedi and Veerabathiran [19, Theorem 3.5]:
@@
where $\rho^{{}^{\prime}}(a)=\big{(}\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}},\phi_{1}(a)\otimes I_{\mathcal{D}_{*,T^{(2)}}}\big{)}$ and $\rho^{{}^{\prime\prime}}(a)=\big{(}\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}},\\ \sigma(a)|_{\mathcal{D}_{*,T^{(2)}}}\big{)}$ are the left actions of $\mathcal{M}$ on the Hilbert spaces $\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}})$ and $(E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}},$ respectively. Indeed, for $h\in\mathcal{H}$ we have
@@
where $U=\begin{bmatrix}A&B\\ C&\widetilde{D}\end{bmatrix}$ defined as above, $\widetilde{D}_{n}=\widetilde{D}(I_{E_{1}}\otimes\widetilde{D})\dots(I_{E_{1}^{\otimes(n-1)}}\otimes\widetilde{D})$ and the series converges with respect to the strong operator topology.
@@ 2.1. Transfer Function
[…]
@@
Thus $\Theta$ is a bi-module map. From [20, Lemma 2.2], $(\rho,S^{(1)},M_{\Theta})$ is a completely contractive covariant representation of $\mathbb{E}$ over $\mathbb{N}_{0}^{2}$ on $\mathcal{F}(E_{1})\otimes\mathcal{D}_{*,T^{(1)}}.$ The contractive linear map $\widetilde{\Theta}$ is called the transfer function of unitary module map $U.$

### v2

@@ Abstract.
The celebrated theorem of Ando says that a pair of commuting contraction on a Hilbert space always dilates to an isometric dilation. Building on this, Solel proved that every completely contractive covariant representation of a product system $\mathbb{E}$ over $\mathbb{N}^{2}_{0}$ dilates to an isometric covariant representation of $\mathbb{E}$ over $\mathbb{N}^{2}_{0}.$ This result generalizes Ando’s theorem in the special case where the $C^{*}$-algebra $\mathcal{M}=\mathbb{C}$ and $\mathbb{E}(\mathbf{n})=\mathbb{C}$ for all $\mathbf{n}\in\mathbb{N}^{2}_{0}.$ This naturally leads to the question: What class of isometric covariant representations of a product system $\mathbb{E}$ over $\mathbb{N}^{2}_{0}$ are sufficient to serve as dilations for completely contractive covariant representations, as established in Solel’s dilation framework? The central objective of this article is to investigate such isometric covariant representations that serve as dilations of completely contractive covariant representations.
@@ 1. Introduction
The starting point of dilation theory is Sz.-Nagy theorem [13], which states that every contraction on a Hilbert space dilates to an isometry on a larger Hilbert space. This result laid the foundation for a deeper understanding of operator theory, providing a powerful tool for studying contractions through their isometric extensions. Ando [1], generalizing Sz.-Nagy’s dilation, constructed isometric dilations for pairs of commuting contractions. That is, for every pair of commuting contractions $(T_{1},T_{2})$ on a Hilbert space $\mathcal{H},$ then there exists a pair of commuting isometries $(V_{1},V_{2})$ on a Hilbert space $\mathcal{K}\supseteq\mathcal{H}$ such that
@@
where $P_{\mathcal{H}}$ denotes the orthogonal projection onto $\mathcal{H}$}. Ando’s theorem confirms that every pair of commuting contractions can be dilated to a pair of commuting isometries, providing a complete solution in the two-variable case. Parrott [14] gave a counterexample showing that for $\geq 3, nn$-tuples of commuting contractions do not generally admit isometric dilations. This striking limitation highlights the intrinsic difficulties of multivariable dilation theory and motivates a more nuanced investigation into the specific conditions under which such dilations might still be possible.
In recent years, a growing body of literature has explored the interplay between dilation and positivity, revealing that certain classes of operator tuples under specific positivity assumptions can indeed possess isometric dilations. See [2, 4, 5, 8, 12] for more details in the polydisc setup. The main motivation behind our interest comes from a recent work by Das and Sarkar [5], obtained an explicit isometric dilation of pair of commuting contractions in the following theorem:
@@
Muhly and Solel [11] presented isometric dilation for completely contractive covariant representations of $C^{*}$-correspondences. Solel [23] proved that every completely contractive covariant representation of product system of $C^{*}$-correspondences over the semigroup $\mathbb{N}_{0}^{2}$ dilates to an isometric representation. Solel [24] studied completely contractive representations of product system of $C^{*}$-correspondences over $\mathbb{N}_{0}^{k},$ and gave a necessary and sufficient condition for such a representation to have a regular isometric dilation. In [21] Skalski established some sufficient conditions for the existence of a not necessarily regular isometric dilation of a completely contractive representation of the product system and discussed the difference between regular and *-regular dilations. The main purpose of this paper is to study such isometric covariant representations that serve as dilations of completely contractive covariant representations.
@@ 1.1. Notations and Preliminaries
We begin this section by recalling few concepts from [9, 11, 15, 18, 20]. Let $E$ be a Hilbert $C^{*}$-module over a $C^{*}$-algebra $\mathcal{M}.$ By $B^{a}(E)$ we denote the $C^{*}$-algebra of all adjointable operators on $E$. We say that the module $E$ is a $C^{*}$-correspondence over $\mathcal{M}$ if there exists a left $\mathcal{M}$-module structure through a non-zero *-homomorphism $\phi:\mathcal{M}\to B^{a}(E)$ in the following sense
@@
for every $\eta_{1},\eta_{2}\in F,$ $\xi_{1},\xi_{2}\in E$ and $a\in\mathcal{M}.$ In this paper, we adopt the following notations: $\mathcal{H}$ denotes a Hilbert space, $E$ represents a $C^{*}$-correspondence over $\mathcal{M}$, and $B(\mathcal{H})$ refers to the algebra of all bounded linear operators on $\mathcal{H}.$ The notion of completely contractive covariant representation plays a crucial role in this paper.
@@
Suppose that $(\sigma,T)$ is a completely bounded covariant representation of E{E} on $\mathcal{H}.$ A closed subspace $\mathcal{W}$ of $\mathcal{H}$ is said to be $(\sigma,T)$-invariant $(resp.(\sigma,T)$-reducing) (cf. [20]) if it is $\sigma(\mathcal{M})$-invariant and (resp. both $\mathcal{W},\mathcal{W}^{\perp}$) is invariant by each operator $T(\xi)$ for all $\xi\in E.$ The restriction provides a new representation $(\sigma,T)|_{\mathcal{W}}$ of $E$ on $\mathcal{W}.$
@@
The following theorem is an analogue of [13, Theorem 3.2].
@@
$\mathcal{H}_{0}:$ $=\{h\in\mathcal{H}:\|(I_{E^{\otimes n}}\otimes\widetilde{T}_{m})(\xi_{n+m}\otimes h)\|=\|\xi_{n+m}\otimes h\|\quad\mbox{and}\quad$
@@ 1.2. Completely contractive representations over product system
Suppose $u=\{u_{1},u_{2},\dots,u_{r}\}\subseteq\{1,2,\dots,k\}$ such that $u_{1}<u_{2}<\dots<u_{r}.$ Define $\mathbf{e}(u):=\mathbf{e}_{u_{1}}+\mathbf{e}_{u_{2}}+\dots+\mathbf{e}_{u_{r}},$ where $\mathbf{e}_{i}:=(0,0,\dots,1,0,\dots 0),$ i.e., $1$ at $i^{th}$ place and 0 otherwise. Next we define the following Brehmer-Solel condition:
@@
The following theorem gave an isometric dilation of completely contractive covariant representations of product system due to Rohilla, Trivedi and Veerabathiran [18, Theorem 3.5]:
@@
where $\rho^{{}^{\prime}}(a)=\big{(}\sigma(a)|_{\mathcal{D}_{*,T^{(1)}}},\phi_{1}(a)\otimes I_{\mathcal{D}_{*,T^{(2)}}}\big{)}$ and $\rho^{{}^{\prime\prime}}(a)=\big{(}\phi_{2}(a)\otimes I_{\mathcal{D}_{*,T^{(1)}}},\sigma(a)|_{\mathcal{D}_{*,T^{(2)}}}\big{)}$ are the left actions of $\mathcal{M}$ on the Hilbert spaces $\mathcal{D}_{*,T^{(1)}}\oplus(E_{1}\otimes\mathcal{D}_{*,T^{(2)}})$ and $(E_{2}\otimes\mathcal{D}_{*,T^{(1)}})\oplus\mathcal{D}_{*,T^{(2)}},$ respectively. Indeed, for $h\in\mathcal{H}$ we have
@@
[…]
@@ 2.1. Transfer Function
where $\xi_{n}\in E_{1}^{\otimes n},\xi\in E_{2},h\in\mathcal{K},n\in\mathbb{N}_{0}$ and $t_{2,1}^{(1,n)}:E_{2}\otimes E_{1}^{\otimes n}\rightarrow E_{1}^{\otimes n}\otimes E_{2}$ is an isomorphism which is a composition of the isomorphisms $\{t_{i,j}\>:\>1\leq i,j\leq 2\}.$ Clearly $M_{\Theta}(\xi)|_{\mathcal{K}}=\Theta(\xi)$ for each $\xi\in E_{2},$ $(\rho,{M}_{\Theta})$ is a c.b.c. representation of $E_{2}$ on $\mathcal{F}(E_{1})\otimes_{\pi}\mathcal{K},$ and it satisfies
@@
Thus $\Theta$ is a bi-module map. From [19, Lemma 2.2], $(\rho,S^{(1)},M_{\Theta})$ is a completely contractive covariant representation of $\mathbb{E}$ over $\mathbb{N}_{0}^{2}$ on $\mathcal{F}(E_{1})\otimes\mathcal{D}_{*,T^{(1)}}.$ The contractive linear map $\widetilde{\Theta}$ is called the transfer function of unitary module map $U.$